OpenStudy (anonymous):

Convert the Cartesian equations to polar form: a. x^2=4y b. x^2-y^2=1 c. y=2x-1 I know the conversion formulas: r=sqroot(x^2+y^2) and theta= tan^-1(y/x) but I keep getting weird answers...I'm in a hurry so please help??!! P.S What I mean by tan^-1 is basically tan-1(u know the inverse of tan)

OpenStudy (tkhunny):

Well, you should do it. I'm not sure I ever would use that inverse tangent substitution. Maybe that's why your answers look funny. Try just \(x = r\cos(\theta)\;and\;y = r\sin(\theta)\)

OpenStudy (p0sitr0n):